1. find f'(x): 1. f(x)=sin(x)+5cos(x) 2. f(x)=(2x+1)(4x+3)

1. find f'(x):

1. f(x)=sin(x)+5cos(x)
2. f(x)=(2x+1)(4x+3)
Answer #1

Solution:

1) f(x)=sin(x)+5cos(x)

\frac{d}{dx}\left(\sin \left(x\right)+5\cos \left(x\right)\right)

\mathrm{Apply\:the\:Sum/Difference\:Rule}:\quad \left(f\pm g\right)^'=f^'\pm g^'

=\frac{d}{dx}\left(\sin \left(x\right)\right)+\frac{d}{dx}\left(5\cos \left(x\right)\right)

=\cos \left(x\right)-5\sin \left(x\right)

2)

f(x)=(2x+1)(4x+3)

\frac{d}{dx}\left(\left(2x+1\right)\left(4x+3\right)\right)

\mathrm{Apply\:the\:Product\:Rule}: (f . g)' = f' . g + f . g'

f=2x+1,\:g=4x+3

=\left(4x+3\right)\frac{d}{dx}\left(2x+1\right)+\left(2x+1\right)\frac{d}{dx}\left(4x+3\right).

\frac{d}{dx}\left(2x+1\right)=2        and    \frac{d}{dx}\left(4x+3\right)=4

=2\left(4x+3\right)+4\left(2x+1\right)

=2\left(8x+5\right)